OpenStudy (anonymous):

The circle formed by the blades of a fan has a circumference of eighty-one and five-sevenths inches. What is its diameter? Use twenty-two sevenths for pi. A. twelve inches B. eighteen inches C.twenty-one inches D.twenty-six inches

OpenStudy (anonymous):

Using the formula for circumference: \[Circumference = \pi \times Diameter\] You know that the circumference is 81 5/7, and pi is 22/7, so: \[81\frac{ 5 }{ 7 } = \frac{ 22 }{ 7 }d\]\[\frac{ 22 }{ 7 }d = \frac{ 572 }{ 7 }\] You can rearrange that equation to find d.

OpenStudy (anonymous):

SO D

OpenStudy (anonymous):

Yep.